Solve and graph the solution set: -2+3(x-3)>=2-(4+x)
Thank you for your time!

Answer by stanbon(75887)

Solve and graph the solution set: -2+3(x-3)>=2-(4+x)
----
-2+3x-9 >= 2-4-x
---
4x-11 >= -2
---
4x >= 9
---
x >=9/4
--------------
Graph:
Draw a number line.
Plot x = 0 and x = 9/4
Draw a solid circle at x = 9/4
Draw an arrow extending to the right from x = 9/4
===============================
Cheers,
Stan H.